OpenStudy (anonymous):

-5f + 40 = 5f +30 f = ?

OpenStudy (anonymous):

-5f + 40 = 5f + 30 40 = 10f + 30 (add 5f to both sides) 10 = 10f (subtract 30 from both sides) F = 1 (divide 10 from both sides)